OpenStudy (anonymous):

exact arc length of y=ln(cos(x)) from x=0 to pi/4 (very important!)

OpenStudy (anonymous):

Well, do you understand how arc length is computed?

OpenStudy (anonymous):

Yes for the most part...this question just has me stuck

OpenStudy (anonymous):

i believe the answer im supposed to get is 1/2ln(3+2sqrt2) but i dont know how

OpenStudy (anonymous):

so get that the derivative would be -sinx/cosx and you plug it into the formula and get sqrt(1+(-sinx/cosx)^2dx from 0 to pi/4 but that is where i am stuck

OpenStudy (freckles):

do you know your pythagorean identities?

OpenStudy (anonymous):

We can parametrize this such that \(x=t\) and \(y=\ln(\cos(t))\). Then we can use \[ds = \sqrt{\left(\frac{dx}{dt}\right)^2 + \left(\frac{dy}{dt}\right)^2}dt\] We are just trying to find: \[ \int ds = \int \sqrt{\left(\frac{dx}{dt}\right)^2 + \left(\frac{dy}{dt}\right)^2}dt \]

OpenStudy (anonymous):

@freckles yes i do know my path identities @wio that is still for arc length?

OpenStudy (freckles):

so you know 1+tan^2(x)=?

OpenStudy (anonymous):

sex^2

OpenStudy (freckles):

yep :)

OpenStudy (anonymous):

sry sec^2 lol

OpenStudy (freckles):

I know what you meant

OpenStudy (anonymous):

For cases where \(y=f(x)\) we let \(x=t\) and our formula simplifies:\[ \int ds = \int \sqrt{1 + \left(\frac{dy}{dx}\right)^2}dx \]

OpenStudy (anonymous):

ohh so since its (-sin^2x/cos^2x) that means its -tan^2? thus being the identity does the negative matter?

OpenStudy (anonymous):

\[ [-\tan(x)]^2 = \tan^2(x) \]

OpenStudy (anonymous):

It should simplify to \(|\sec x|\).

OpenStudy (anonymous):

ohh lmfao didnt cath that thanks therefore i get sqrtsecx^2 then giving me the absolute value

OpenStudy (anonymous):

However, on the \((0,\pi/4)\) interval, I believe that \(cos\) remains positive, so \(\sec\) should do so as well.

OpenStudy (anonymous):

k so then i just do the anti derivative and plug in 0 and pi/4 correct?

OpenStudy (alekos):

I don't think the answer is 1/2ln(3+2sqrt2)

OpenStudy (alekos):

I get \[\ln (1+\sqrt{2})\]

OpenStudy (anonymous):

yes thank you i appreciate the help @alekos after he @wio and @freckles explained the steps i got the same answe...Wish i could at least give them both medals but i cannot i gresstly appreciate everyones help :D